Could you please help me solve this problem; A train leaves San Fran for L.A. at 8:00am. A second train leaves L.A. for San Fran at 8:00am the same day. The distance between the two cities is 400 miles. If each train travels at an average speed of 80 miles per hour, what time will it be when the two trains meet?
Found 2 solutions by solver91311, xinxin:
Answer by solver91311(24713)



How long does it take to go 400 miles at 160 miles per hour? Add that amount of time to 8:00AM. Hint: distance equals rate times time, so time equals distance divided by rate.

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by xinxin(76)

Suppose the route from San Fran to L.A is a straight line. If the two trains travel at the same speed and depart at the same time, they should meet at the center point of this line. The total distance is 400 miles, so half of it is 200 miles.
The time it takes to travel 200 miles = 200/80 = 2.5 hours, thus the two will meet at 10:30 AM (= 8 + 2.5).